### v1

# AGE-Net: Spectral–Spatial Fusion and Anatomical Graph Reasoning with Evidential Ordinal Regression for Knee Osteoarthritis Grading

Abstract
Automated Kellgren–Lawrence (KL) grading from knee radiographs is challenging due to subtle structural changes, long-range anatomical dependencies, and ambiguity near grade boundaries. We propose AGE-Net, a ConvNeXt-based framework integrating Spectral–Spatial Fusion (SSF), Anatomical Graph Reasoning (AGR), and Differential Refinement (DFR). To capture predictive uncertainty and preserve label ordinality, AGE-Net further employs a Normal-Inverse-Gamma (NIG) evidential regression head (COE-Head) and a pairwise ordinal ranking constraint. On a knee KL dataset, AGE-Net achieves QWK 0.90170.0045 and MSE 0.23490.0028 (3 seeds), improving over strong CNN baselines and confirming consistent gains in ablation studies. We additionally outline uncertainty quality, robustness, and explainability analyses and reserve space for upcoming experimental figures.
Keywords: Knee osteoarthritis; Kellgren–Lawrence grading; Ordinal regression; Graph reasoning; Evidential learning; Uncertainty
1 Introduction
Knee osteoarthritis (OA) is among the most prevalent degenerative musculoskeletal disorders, imposing substantial burden through chronic pain, reduced mobility, and diminished quality of life. Radiographic Kellgren–Lawrence (KL) grading Kellgren and Lawrence (1957) remains one of the most widely adopted clinical standards to characterize OA severity, typically ranging from Grade 0 (no OA) to Grade 4 (severe OA). Automating KL grading from plain radiographs could reduce workload and inter-observer variation and facilitate population-scale screening.
Despite progress in deep learning for radiography, reliable KL grading remains difficult due to three recurring challenges. First, KL transitions may be driven by subtle and localized findings, such as early osteophytes or mild joint space narrowing, frequently embedded in low-contrast anatomical contexts. Second, OA severity is influenced by long-range anatomical structure: medial and lateral compartments, tibiofemoral alignment, and spatially distributed textural cues jointly contribute to an overall grade, which can be under-modeled by strictly local operations. Third, KL labels are ordinal and inherently uncertain near boundaries; borderline cases are common, and label noise is non-negligible even among expert readers. A clinically useful system should thus (i) preserve monotonic ordering across grades and (ii) represent uncertainty that correlates with ambiguity and error.
Modern CNN backbones (e.g., ConvNeXt Liu et al. (2022)) provide strong feature extraction, but end-to-end pipelines often treat KL as plain classification and do not explicitly model anatomy-aligned long-range dependencies, nor do they provide calibrated uncertainty. We propose AGE-Net, a ConvNeXt-based framework with four key components: (1) Spectral–Spatial Fusion (SSF) for frequency-sensitive and spatially recalibrated enhancement; (2) Anatomical Graph Reasoner (AGR) to model long-range anatomical relations by graph message passing on pooled tokens; (3) Differential Refiner (DFR) to emphasize edge- and boundary-sensitive cues that often indicate joint degeneration; and (4) COE-Head evidential NIG regression Amini et al. (2020) plus a pairwise ordinal ranking constraint to enforce ordinal consistency and support uncertainty-aware decision making.
Contributions. This paper makes the following contributions:
We introduce AGE-Net, a unified KL grading framework integrating SSF, AGR, and DFR on top of a strong ConvNeXt backbone Liu et al. (2022).
We incorporate evidential NIG regression Amini et al. (2020) and a pairwise ordinal ranking constraint to improve uncertainty modeling and ordinal consistency.
We report strong empirical performance with ablation evidence and reserve additional experiments (uncertainty quality, robustness, explainability) with dedicated figure placeholders for ongoing work.
2 Related Work
2.1 Backbones for radiographs and medical imaging
CNN backbones such as VGG Simonyan and Zisserman (2015), ResNet He et al. (2016), DenseNet Huang et al. (2017), and Inception Szegedy et al. (2016) are common choices in radiograph analysis. EfficientNet Tan and Le (2019) improves scaling and parameter efficiency. ConvNeXt Liu et al. (2022) modernizes ConvNet design and offers strong transfer learning performance. Vision Transformers such as ViT Dosovitskiy et al. (2021) and Swin Liu et al. (2021) provide global modeling, but can be sensitive to pretraining and optimization, especially in medical domains with limited labeled data and distribution shifts.
2.2 Ordinal regression and ranking constraints
KL grades are ordinal; treating them as independent classes can yield inconsistent decision boundaries and suboptimal calibration around adjacent grades. Ordinal learning frameworks, including rank-consistent formulations like CORAL Cao et al. (2020), aim to preserve monotonic structure. Pairwise ranking constraints provide a practical mechanism to enforce ordering by penalizing violations on valid grade-separated pairs.
2.3 Graph reasoning for non-local dependencies
Graph neural networks (GNNs) and related message passing mechanisms can explicitly model non-local dependencies among tokens or anatomical regions. EdgeConv and dynamic graph construction ideas (e.g., DGCNN) Wang et al. (2019) motivate our anatomy-aware graph reasoning over pooled radiographic tokens.
2.4 Uncertainty estimation and evidential learning
Uncertainty estimation is critical for borderline and ambiguous KL cases. Evidential learning Amini et al. (2020) enables deterministic prediction of distributional parameters without sampling-heavy Bayesian inference, making it attractive for practical deployment and uncertainty-aware triage.
3 Method
3.1 Problem formulation
We consider ordinal KL grading with labels . Instead of treating grades as independent classes, we produce a continuous score and map to grade by rounding/clipping during evaluation. The model should (i) predict accurate scores and (ii) reflect uncertainty associated with ambiguous cases.
3.2 Backbone and feature map
Given an input radiograph , the ConvNeXt backbone Liu et al. (2022) outputs a deep feature map
AGE-Net augments with three enhancement modules (SSF, AGR, DFR), followed by global pooling and a prediction head.
3.3 Spectral–Spatial Fusion (SSF)
Motivation. Radiographic OA cues often manifest as texture variations and subtle contrast changes. Convolutional features can underutilize global frequency context. SSF introduces lightweight global frequency mixing and spatial recalibration, encouraging the model to capture both textural patterns (frequency domain) and anatomy-aligned emphasis (spatial domain).
Spectral path. Let denote the 2D Fourier transform (implemented via rFFT for real signals). We compute
We apply channel-wise learnable complex scaling :
This performs a global, channel-wise frequency reweighting that can boost frequency components linked to fine-grained radiographic texture.
Spatial attention path. We compute a spatial gate , where is a lightweight bottleneck:
with a nonlinearity (e.g., ReLU), and sigmoid.
Fusion. The fused output uses a residual stabilization:
Residual addition mitigates optimization instability and preserves backbone semantics, while enabling SSF to act as a learnable enhancement.
Complexity. SSF adds negligible parameters (complex scale per channel plus two conv layers). FFT cost is but remains practical at the feature-map resolution (post-backbone) and is amortized across training.
3.4 Anatomical Graph Reasoner (AGR)
Motivation. KL grading depends on relationships across anatomical compartments. Standard convolutions may fail to integrate long-range structure efficiently. AGR constructs an anatomy-aware token graph from pooled features and performs message passing to capture non-local dependencies.
Tokenization. We reduce channels and pool to a fixed grid:
Flatten spatially to obtain node features:
kNN graph construction. We build a kNN graph in feature space (k=9). For each node , neighbors are selected by nearest distances, creating an adaptive topology that can align with anatomical variability.
EdgeConv-style aggregation. Inspired by DGCNN Wang et al. (2019), we compute edge features
where is a shared MLP (implemented as conv + BN + LeakyReLU). Aggregation uses max pooling:
We reshape back to and upsample to , then map to a gate :
Gated residual. AGR returns
The gate emphasizes anatomy-aligned features while preserving base representation.
Complexity. Graph building costs for naive pairwise distances; with this is manageable. Message passing scales with and is lightweight compared with backbone compute.
3.5 Differential Refiner (DFR)
Motivation. OA severity is often indicated by boundary and edge changes (e.g., joint space narrowing contours). DFR explicitly highlights differential responses between a depthwise-filtered feature and the original feature map, encouraging sensitivity to subtle structural transitions.
Depthwise response and differential. We compute a depthwise response :
and define an edge-like differential map
Channel mixing and stabilization. We apply a mixing conv and normalization:
Finally, DFR outputs a residual nonlinearity:
Interpretation. The absolute difference acts like a learned gradient operator in feature space, while the residual ensures that DFR acts as refinement rather than a destructive transform.
Complexity. DFR adds one depthwise conv and one pointwise conv per block; parameter overhead is modest and suitable for high-resolution inputs.
3.6 COE-Head: Evidential NIG Regression
Motivation. Borderline KL cases are intrinsically ambiguous; a model should not only output a score, but also express uncertainty. Evidential regression Amini et al. (2020) predicts distributional parameters that capture both aleatoric and epistemic uncertainty signals in a deterministic forward pass.
Parameterization. The COE head maps pooled embedding to four parameters:
where is the predictive mean, and define the Normal-Inverse-Gamma prior. We enforce positivity constraints via softplus:
Predictive uncertainty. Under NIG, the induced predictive distribution is a Student-; a commonly used predictive variance is
Higher evidence (large ) typically yields narrower predictive distributions.
Loss. We optimize NIG negative log-likelihood plus evidence regularization:
with warm-up schedule matching the current training code.
3.7 Ordinal Ranking Constraint
Motivation. KL grades encode monotonic severity; ordinal consistency is important to avoid implausible outputs (e.g., predicting lower score for obviously more severe images). Pairwise constraints provide a direct mechanism to enforce ordering.
Valid pairs and margin. For a mini-batch, define the set of valid pairs
We enforce a margin on scores:
We disable ranking under Mixup because labels become soft and the pairwise inequality becomes ill-defined, consistent with the current implementation.
Total objective. The final training objective is
with in our experiments.
3.8 End-to-end pipeline
Let and apply modules sequentially:
We pool to embedding , predict NIG parameters via COE, and compute total loss . In ablations, we disable each module by removing its transform from the sequence.
4 Experiments
4.1 Dataset and evaluation protocol
We evaluate on a knee radiograph dataset with KL labels . We repeat all runs with 3 random seeds and report meanstd. (TBD) We will document dataset size, acquisition protocol, per-grade distribution, and patient-level split strategy to avoid subject leakage across train/val/test sets.
4.2 Implementation details
Inputs are resized to and randomly cropped to . We train for 300 epochs using AdamW Loshchilov and Hutter (2019) with cosine annealing Loshchilov and Hutter (2017). We apply RandomAffine and ColorJitter augmentations, Mixup Zhang et al. (2018) with probability 0.5, and Random Erasing Zhong et al. (2020) with probability 0.25. Training uses mixed precision and EMA (decay 0.999). Test-time augmentation (TTA) averages predictions of an image and its horizontal flip.
| Item | Setting |
| Backbone | ConvNeXt-Base Liu et al. (2022) (offline weights) |
| Input size | |
| Epochs | 300 |
| Optimizer | AdamW Loshchilov and Hutter (2019) |
| Base LR | (head), backbone |
| Schedule | Cosine annealing Loshchilov and Hutter (2017) |
| Weight decay | |
| EMA | decay 0.999 |
| Mixup | Zhang et al. (2018) |
| Random Erasing | Zhong et al. (2020) |
| TTA | flip average |
| Ranking | , |
4.3 Metrics
We report quadratic weighted kappa (QWK) Cohen (1968), mean squared error (MSE), accuracy (ACC), macro-F1, and recall. For QWK/ACC/F1/Recall, continuous predictions are rounded to nearest integer grade and clipped to .
5 Results
5.1 Main comparison
Table 2 reports that AGE-Net achieves the best overall performance (QWK 0.90170.0045; MSE 0.23490.0028) compared with strong CNN baselines and representative transformer baselines under our current setting.
| Method | QWK | MSE | ACC | F1 | Rec |
| AGE-Net (Full) | 0.90170.0045 | 0.23490.0028 | 0.78390.0056 | 0.80120.0053 | 0.79190.0060 |
| AGE-Net (w/o Rank) | 0.89630.0033 | 0.26000.0083 | 0.77380.0056 | 0.78930.0067 | 0.77970.0056 |
| AGE-Net (w/o AGR) | 0.89320.0018 | 0.26650.0017 | 0.76820.0107 | 0.77860.0131 | 0.76340.0149 |
| AGE-Net (Base ConvNeXt) | 0.89090.0113 | 0.27950.0287 | 0.77010.0051 | 0.77740.0273 | 0.76320.0373 |
| VGG16 Simonyan and Zisserman (2015) | 0.86540.0069 | 0.30460.0102 | 0.71540.0089 | 0.71800.0170 | 0.69550.0165 |
| EfficientNet Tan and Le (2019) | 0.86380.0012 | 0.30780.0060 | 0.70870.0049 | 0.71660.0050 | 0.69390.0058 |
| ConvNeXt-Base Liu et al. (2022) | 0.86020.0077 | 0.32480.0118 | 0.71320.0087 | 0.72950.0019 | 0.72980.0039 |
| Inception-V3 Szegedy et al. (2016) | 0.84750.0075 | 0.32520.0069 | 0.67520.0030 | 0.68550.0093 | 0.65900.0113 |
| DenseNet121 Huang et al. (2017) | 0.84270.0088 | 0.32740.0032 | 0.66290.0126 | 0.60100.0191 | 0.57500.0189 |
| ResNet50 He et al. (2016) | 0.82710.0012 | 0.37110.0021 | 0.62720.0033 | 0.59320.0251 | 0.56290.0195 |
| Swin-T Liu et al. (2021) | 0.56030.0329 | 0.77670.0493 | 0.35310.0140 | 0.29350.0179 | 0.33630.0101 |
| ViT-B Dosovitskiy et al. (2021) | 0.34730.0298 | 1.04130.0084 | 0.26820.0168 | 0.16120.0217 | 0.25190.0112 |
5.2 Ablation study
Table 3 isolates the effect of major components. Removing ranking decreases QWK, consistent with the importance of explicit ordinal constraints. Removing AGR leads to a larger drop, highlighting the value of anatomy-aware long-range reasoning.
| Variant | QWK | MSE |
| Full | 0.90170.0045 | 0.23490.0028 |
| w/o Rank | 0.89630.0033 | 0.26000.0083 |
| w/o AGR | 0.89320.0018 | 0.26650.0017 |
| Base ConvNeXt | 0.89090.0113 | 0.27950.0287 |
| (TBD) w/o SSF | – | – |
| (TBD) w/o DFR | – | – |
| (TBD) w/o COE (heteroscedastic) | – | – |
5.3 Per-grade analysis (planned / in-progress)
We will add per-grade precision/recall and confusion matrices to better characterize borderline failure modes (e.g., confusion between grades 1/2/3). These analyses often reveal whether improvements concentrate on adjacent grade disambiguation.
5.4 Uncertainty quality (planned / in-progress)
To validate uncertainty, we will evaluate: (i) binned uncertainty vs. absolute error; (ii) selective prediction via risk–coverage curves; (iii) correlation metrics (Spearman/Pearson) between uncertainty and error.
5.5 Robustness under perturbations (planned / in-progress)
We will evaluate robustness under brightness/contrast shifts, Gaussian noise, blur, and occlusion. We will report QWK/MSE degradation curves and compare baseline vs. AGE-Net to quantify resilience benefits from SSF/DFR/AGR.
6 Discussion
6.1 Why SSF helps
SSF provides global frequency mixing and spatial recalibration. In radiographs, fine texture patterns and subtle contrast differences can be critical for early OA cues. Frequency-domain modulation enables the model to learn global spectral emphasis, while spatial gating focuses enhancement on informative regions. The residual design stabilizes optimization and keeps SSF lightweight.
6.2 Why AGR helps
AGR explicitly introduces non-local reasoning over pooled tokens. This is particularly relevant for tibiofemoral joint assessment, where cues span multiple compartments. By constructing kNN graphs in feature space, AGR can adapt its neighborhood structure to anatomical variability and encode long-range relations beyond fixed receptive fields. The gating mechanism provides an interpretable pathway for anatomy-aligned emphasis.
6.3 Why DFR helps
DFR targets edge- and boundary-sensitive cues. KL grading relies on joint space narrowing and contour changes; feature differentials can highlight precisely such transitions. Unlike heavy edge detectors in pixel space, DFR learns edge sensitivity in feature space and integrates seamlessly with residual learning.
6.4 Uncertainty and ordinality
COE evidential regression provides a mechanism to represent uncertainty that can be used for triage (e.g., defer high-uncertainty cases). The ranking constraint enforces monotonicity and combats degenerate regressors that might fit MSE while violating grade ordering. Together, they align the predictor with clinical expectations: ordinal consistency and uncertainty awareness.
6.5 Limitations and future work
First, we will complete uncertainty calibration, robustness testing, and explainability analyses and include the corresponding figures. Second, transformer baselines may require stronger pretraining or alternative optimization; we will add tuned versions and report results. Third, generalization across acquisition sites and demographic subgroups will be studied with external validation if available.
7 Conclusion
We presented AGE-Net for ordinal KL grading from knee radiographs. AGE-Net integrates SSF, AGR, and DFR with a COE evidential NIG head and an ordinal ranking constraint. The method achieves strong performance (QWK 0.90170.0045; MSE 0.23490.0028) and shows consistent improvements in ablations. Ongoing experiments will further quantify uncertainty quality, robustness, and interpretability.
Appendix A Additional implementation notes
A.1 Training tricks and stability
We adopt mixed precision training to improve throughput, and EMA to stabilize evaluation. Layer-wise learning rate decouples backbone and head updates: the backbone uses a smaller learning rate to reduce catastrophic forgetting of pretrained features, while the head learns task-specific mapping. We disable ranking loss under Mixup because labels become interpolated and the ordering constraint becomes ill-defined in soft label space.
A.2 Test-time augmentation
We apply flip-based TTA by averaging predictions of the original and horizontally flipped inputs. This can reduce variance and improve robustness, especially for noisy radiographs.
Appendix B Planned extended ablations
w/o SSF: remove frequency modulation and spatial gating to isolate spectral contribution.
w/o DFR: remove differential refinement to isolate boundary-sensitive enhancement.
w/o COE: replace evidential head with heteroscedastic regression head and compare uncertainty quality.
k sensitivity: vary k in AGR (e.g., 5/9/13) to study neighborhood size effects.
Token grid size: vary pooling grid (e.g., 1010, 1414, 1616).
Appendix C Statistical significance (planned)
We will report paired significance tests across seeds (e.g., paired -tests on per-seed QWK) and effect sizes when comparing full AGE-Net with major ablations and the ConvNeXt baseline.