### v2

# AGE-Net: Spectral–Spatial Fusion and Anatomical Graph Reasoning with Evidential Ordinal Regression for Knee Osteoarthritis Grading

Abstract
Purpose: To develop and evaluate an ordinal, uncertainty-aware deep learning framework for automated Kellgren–Lawrence (KL) grading from knee radiographs, addressing subtle degenerative cues, long-range anatomical dependencies, and ambiguity near adjacent-grade boundaries.
Materials and Methods: In this retrospective study, we proposed AGE-Net, built on a ConvNeXt-Base backbone and augmented with (i) Spectral–Spatial Fusion (SSF) to enhance frequency-sensitive textures and spatially reweight informative regions, (ii) Anatomical Graph Reasoning (AGR) to model non-local anatomy-aligned dependencies, and (iii) Differential Refinement (DFR) to amplify boundary-sensitive degeneration cues. To preserve label ordinality and quantify uncertainty, AGE-Net employed a Normal–Inverse-Gamma (NIG) evidential regression head (COE-Head) together with a pairwise ordinal ranking constraint. Model performance was evaluated on a knee radiograph KL dataset (https://www.kaggle.com/datasets/shashwatwork/knee-osteoarthritis-dataset-with-severity) using three random seeds; results are reported as mean SD. Primary metrics were quadratic weighted kappa (QWK) and mean squared error (MSE).
Results: Across three random seeds, AGE-Net achieved a QWK of and an MSE of , outperforming strong convolutional baselines under the same setting. Ablation experiments demonstrated complementary gains from SSF, AGR, DFR, and the ordinal–evidential design.
Conclusion: An ordinal, uncertainty-aware architecture that integrates frequency-enhanced feature extraction and anatomy-aligned non-local reasoning improved automated KL grading on knee radiographs, supporting its potential utility for scalable osteoarthritis severity assessment.
1 Introduction
Knee osteoarthritis (OA) is among the most prevalent degenerative musculoskeletal disorders, imposing substantial burden through chronic pain, reduced mobility, and diminished quality of life. Radiographic Kellgren–Lawrence (KL) grading Kellgren and Lawrence (1957) remains one of the most widely adopted clinical standards to characterize OA severity, typically ranging from Grade 0 (no OA) to Grade 4 (severe OA). Automating KL grading from plain radiographs could reduce workload and inter-observer variation and facilitate population-scale screening.
Despite progress in deep learning for radiography, reliable KL grading remains difficult due to three recurring challenges. First, KL transitions may be driven by subtle and localized findings, such as early osteophytes or mild joint space narrowing, frequently embedded in low-contrast anatomical contexts. Second, OA severity is influenced by long-range anatomical structure: medial and lateral compartments, tibiofemoral alignment, and spatially distributed textural cues jointly contribute to an overall grade, which can be under-modeled by strictly local operations. Third, KL labels are ordinal and inherently uncertain near boundaries; borderline cases are common, and label noise is non-negligible even among expert readers. A clinically useful system should thus (i) preserve monotonic ordering across grades and (ii) represent uncertainty that correlates with ambiguity and error.
Modern CNN backbones (e.g., ConvNeXt Liu et al. (2022)) provide strong feature extraction, but end-to-end pipelines often treat KL as plain classification and do not explicitly model anatomy-aligned long-range dependencies, nor do they provide calibrated uncertainty. We propose AGE-Net, a ConvNeXt-based framework with four key components: (1) Spectral–Spatial Fusion (SSF) for frequency-sensitive and spatially recalibrated enhancement; (2) Anatomical Graph Reasoner (AGR) to model long-range anatomical relations by graph message passing on pooled tokens; (3) Differential Refiner (DFR) to emphasize edge- and boundary-sensitive cues that often indicate joint degeneration; and (4) COE-Head evidential NIG regression Amini et al. (2020) plus a pairwise ordinal ranking constraint to enforce ordinal consistency and support uncertainty-aware decision making.
Contributions. This paper makes the following contributions:
We propose AGE-Net, a unified framework for automated Kellgren–Lawrence (KL) grading that integrates Spectral–Spatial Fusion (SSF), Anatomical Graph Reasoning (AGR), and Differential Refinement (DFR) on top of a ConvNeXt backbone Liu et al. (2022) to capture frequency-aware textures, anatomy-aligned non-local dependencies, and boundary-sensitive degenerative cues.
We introduce an evidential ordinal regression formulation by coupling Normal–Inverse-Gamma (NIG) evidential regression Amini et al. (2020) with a pairwise ordinal ranking constraint, improving ordinal consistency while enabling uncertainty-aware assessment for ambiguous cases.
We provide comprehensive empirical validation, including comparisons to strong baselines and component-wise ablations that quantify the contribution of each module under a consistent evaluation protocol.
2 Related Work
2.1 Backbones for radiographs and medical imaging
Convolutional neural networks (CNNs) remain the dominant backbone family for radiograph analysis due to their strong inductive bias and stable optimization on modest-sized medical datasets. Classic architectures such as VGG Simonyan and Zisserman (2015), ResNet He et al. (2016), DenseNet Huang et al. (2017), and Inception Szegedy et al. (2016) have been widely adopted, while EfficientNet Tan and Le (2019) further improves accuracy–efficiency trade-offs through principled compound scaling. More recently, ConvNeXt Liu et al. (2022) revisits and modernizes ConvNet design, yielding strong transfer performance with training recipes comparable to Transformer-based models. In parallel, Vision Transformers (ViT) Dosovitskiy et al. (2021) and hierarchical variants such as Swin Liu et al. (2021) enable global context modeling; however, in medical imaging they often exhibit higher sensitivity to pretraining, hyperparameters, and distribution shift, particularly when labeled data are limited. These considerations motivate our choice of a modern CNN backbone (ConvNeXt) as a robust foundation for KL grading, while introducing explicit modules to capture non-local anatomy-aligned dependencies and ordinal ambiguity.
2.2 Ordinal regression and ranking constraints
KL grading is inherently ordinal: errors are not equally harmful (e.g., predicting Grade 4 as Grade 3 is less severe than as Grade 0), and adjacent grades often differ only by subtle radiographic cues. Consequently, casting KL grading as a standard multi-class classification problem can produce decision boundaries that ignore monotonicity, yield label-incoherent predictions across thresholds, and behave unstably in boundary cases where inter-reader variability is common.
A typical ordinal learning strategy reformulates a -grade problem into ordered binary subproblems with shared representation, such that the predicted probabilities are rank-consistent across thresholds. CORAL Cao et al. (2020), for example, enforces consistent ordering by learning a single latent score with multiple ordered cutpoints (or equivalently, multiple binary heads with constrained weights), thereby encouraging monotone cumulative predictions. Alternative approaches treat ordinality via regression to a continuous severity score followed by discretization; while simple, they may underutilize the discrete threshold structure and can be sensitive to class imbalance and heteroscedastic noise.
Complementary to threshold-based formulations, pairwise ranking constraints enforce ordering at the sample level. Given two radiographs and with , a ranking loss penalizes violations of the desired order by encouraging the model score to satisfy , where is a margin. In practice, the loss can be implemented with a hinge form,
or a smooth logistic alternative. Such constraints are attractive because they (i) directly encode the monotonic prior, (ii) can be selectively applied to valid grade-separated pairs (e.g., adjacent-only or ), and (iii) mitigate boundary ambiguity by emphasizing relative ordering rather than absolute class identity. These properties make pairwise ranking a practical add-on to ordinal regression for KL grading, especially under noisy labels and imbalanced grade distributions.
2.3 Graph reasoning for non-local dependencies
Non-local dependencies are pervasive in musculoskeletal radiographs: clinically relevant cues (e.g., joint-space narrowing, osteophytes, and subchondral sclerosis) are spatially distributed and must be interpreted relative to anatomical context, such as tibiofemoral alignment and medial–lateral compartment configuration. Graph neural networks (GNNs) provide a convenient inductive bias for such settings by representing regions (or tokens) as nodes and propagating information through message passing, enabling structured aggregation beyond local receptive fields.
Early message passing formulations and modern GNN variants differ mainly in how neighborhoods are defined and how edge-specific interactions are parameterized. In computer vision, dynamic graph construction has been shown effective for capturing long-range and content-adaptive relationships. For example, DGCNN introduces EdgeConv and recomputes graph neighborhoods in feature space, allowing the connectivity to evolve with the representation and encouraging appearance-aware interactions Wang et al. (2019). Similar principles have been adopted in graph-based reasoning modules that operate on compact sets of region descriptors, where each node summarizes a local patch while edges encode similarity or proximity.
In radiograph analysis, graph reasoning is particularly appealing because it can: (i) aggregate evidence across anatomically correlated regions (e.g., femoral condyle and tibial plateau), (ii) capture asymmetric compartmental patterns (medial vs. lateral), and (iii) improve robustness to scale and position variations by reasoning over relational structure rather than absolute pixel locations. These motivations underpin our anatomy-aware graph reasoning design, which performs message passing over a set of pooled radiographic tokens to explicitly model non-local anatomical dependencies while keeping computation tractable.
Relation to self-attention. Compared with global self-attention, which forms dense all-to-all interactions and can be data- and regularization-sensitive in medical settings, graph reasoning provides an explicit sparse relational prior through controllable neighborhoods. This structure can reduce spurious long-range mixing, improve interpretability at the level of node/edge interactions, and offer a computationally efficient mechanism to capture anatomically meaningful non-local dependencies, making it complementary to attention-based modeling.
2.4 Uncertainty estimation and evidential learning
Uncertainty estimation is particularly important for KL grading because borderline cases are common and adjacent grades can be difficult to distinguish even for expert readers. In this context, it is useful to distinguish aleatoric uncertainty, which arises from intrinsic ambiguity and label noise (e.g., subtle osteophytes or equivocal joint-space narrowing), from epistemic uncertainty, which reflects limited knowledge of the model (e.g., out-of-distribution anatomy, acquisition differences, or rare presentations). Reliable uncertainty estimates can support uncertainty-aware triage (e.g., flagging low-confidence cases for secondary review) and reduce the risk of overconfident misgrading.
Common uncertainty approaches include Bayesian neural networks and approximate Bayesian inference, Monte Carlo dropout, deep ensembles, and post-hoc calibration techniques; however, many of these require multiple stochastic forward passes, sampling, or maintaining several models, increasing computational cost and complicating deployment. Evidential learning offers an attractive alternative by producing uncertainty estimates deterministically in a single forward pass. In particular, the Normal–Inverse-Gamma (NIG) evidential regression framework Amini et al. (2020) predicts the parameters of a predictive distribution together with an evidence quantity that reflects the model’s confidence, thereby capturing input-dependent uncertainty without sampling-heavy Bayesian inference.
Beyond efficiency, evidential formulations provide a principled way to couple prediction accuracy with uncertainty via evidence regularization: when the model is wrong or the data are ambiguous, it is encouraged to reduce evidence (increase uncertainty), while confident predictions are supported only when justified by the data. These properties make evidential learning well suited to KL grading, where uncertainty is clinically meaningful and often concentrated around adjacent-grade boundaries.
3 Method
3.1 Task definition and outputs
We study automated Kellgren–Lawrence (KL) grading from knee radiographs, where the reference label is an ordinal grade . Rather than modeling KL as five independent classes, we learn a continuous severity score and convert it to a discrete grade by rounding and clipping at evaluation time. This formulation reflects the ordinal nature of KL and supports uncertainty-aware assessment for borderline cases. Accordingly, the model is designed to (i) predict accurate severity scores and (ii) estimate predictive uncertainty that increases for ambiguous inputs and decreases when evidence is strong.
3.2 Overview of AGE-Net
AGE-Net uses a modern convolutional backbone to extract a deep feature map from an input radiograph, then applies three lightweight enhancement modules that target complementary failure modes in KL grading: (i) Spectral–Spatial Fusion (SSF) for frequency-aware texture enhancement and spatial recalibration, (ii) Anatomical Graph Reasoning (AGR) for anatomy-aligned non-local dependency modeling, and (iii) Differential Refinement (DFR) to emphasize boundary-sensitive structural changes. A global pooling operator produces an image-level embedding, which is mapped to an evidential regression head (COE-Head) that predicts a severity score and its uncertainty. An additional pairwise ordinal ranking constraint is used during training to encourage monotonic ordering across grades.
3.3 Backbone feature extraction
Given an input radiograph , the ConvNeXt backbone Liu et al. (2022) outputs a deep feature map
The subsequent modules operate on and preserve its spatial resolution, enabling localized degenerative cues to be enhanced while maintaining global anatomical context.
3.4 Spectral–Spatial Fusion (SSF)
Rationale. Radiographic OA findings may present as subtle, spatially localized texture and contrast changes. To better leverage global frequency context while retaining spatial selectivity, SSF combines frequency-domain modulation with a lightweight spatial gate.
Spectral modulation. Let denote the 2D Fourier transform (implemented via rFFT for real-valued features). We compute
and apply channel-wise learnable complex scaling :
This performs global, channel-wise frequency reweighting to emphasize components associated with fine-grained radiographic patterns.
Spatial recalibration. We compute a spatial gate using a lightweight bottleneck:
where is a nonlinearity (e.g., ReLU) and is the sigmoid.
Fusion. SSF outputs a residual-stabilized enhancement:
Computation. SSF introduces minimal additional parameters (complex scale per channel and two convolutions). FFT cost is at post-backbone resolution.
3.5 Anatomical Graph Reasoning (AGR)
Rationale. KL grading depends on relationships across anatomical compartments and non-local structures. AGR introduces explicit message passing over pooled tokens to capture anatomy-aligned non-local dependencies in a computationally controlled manner.
Tokenization. We reduce channel dimensionality and pool to a fixed grid:
Flattening yields node features
kNN graph construction. A kNN graph is constructed in feature space with ; for each node , neighbors are selected by nearest distances. This adaptive topology accommodates anatomical variability across radiographs.
EdgeConv-style message passing. Inspired by DGCNN Wang et al. (2019), we compute edge features
where is a shared MLP (implemented as conv + BN + LeakyReLU). Aggregation uses max pooling:
We reshape back to , upsample to , and map to a sigmoid gate:
Gated residual update.
Computation. With tokens, neighborhood search and message passing are lightweight relative to backbone computation.
3.6 Differential Refinement (DFR)
Rationale. OA severity can be expressed through boundary and contour changes (e.g., joint-space narrowing). DFR highlights differential responses between a depthwise-filtered feature map and the original map to emphasize boundary-sensitive cues.
Differential map. We compute a depthwise response
and an absolute differential map
Refinement. A pointwise mixing layer and normalization produce
and DFR returns a residual nonlinearity
3.7 Evidential regression head (COE-Head)
Rationale. Borderline KL cases are intrinsically ambiguous; therefore, the model should output both a severity score and an uncertainty estimate. We adopt evidential regression Amini et al. (2020), which predicts distributional parameters deterministically (single forward pass), making it suitable for uncertainty-aware triage.
Parameterization. Let denote the pooled embedding. The COE-Head predicts
where is the predictive mean and parameterize a Normal–Inverse-Gamma prior. Positivity constraints are enforced via
Predictive uncertainty. The induced predictive distribution is Student-; a commonly used predictive variance is
Loss. We optimize the evidential objective comprising NIG negative log-likelihood and evidence regularization:
with the warm-up schedule .
3.8 Ordinal ranking constraint
Rationale. To encourage ordinal consistency, we apply a pairwise ranking constraint that penalizes violations of monotonic ordering.
Valid pairs and margin. Within a mini-batch, we define valid grade-separated pairs
With margin , the ranking loss is
Ranking is disabled under Mixup augmentation because labels become soft and the pairwise inequality is no longer well-defined.
Total objective.
where .
3.9 End-to-end inference pipeline
Let and apply modules sequentially:
We pool to obtain , predict using the COE-Head, and optimize . At inference, the predicted continuous score is mapped to KL grade by rounding and clipping to . In ablation studies, each module is removed from the sequence while keeping the remaining components unchanged.
4 Experiments
4.1 Dataset and evaluation protocol
We evaluate on a knee radiograph dataset with KL labels . We repeat all runs with 3 random seeds and report meanstd. (TBD) We will document dataset size, acquisition protocol, per-grade distribution, and patient-level split strategy to avoid subject leakage across train/val/test sets.
4.2 Implementation details
Inputs are resized to and randomly cropped to . We train for 300 epochs using AdamW Loshchilov and Hutter (2019) with cosine annealing Loshchilov and Hutter (2017). We apply RandomAffine and ColorJitter augmentations, Mixup Zhang et al. (2018) with probability 0.5, and Random Erasing Zhong et al. (2020) with probability 0.25. Training uses mixed precision and EMA (decay 0.999). Test-time augmentation (TTA) averages predictions of an image and its horizontal flip.
| Item | Setting |
| Backbone | ConvNeXt-Base Liu et al. (2022) (offline weights) |
| Input size | |
| Epochs | 300 |
| Optimizer | AdamW Loshchilov and Hutter (2019) |
| Base LR | (head), backbone |
| Schedule | Cosine annealing Loshchilov and Hutter (2017) |
| Weight decay | |
| EMA | decay 0.999 |
| Mixup | Zhang et al. (2018) |
| Random Erasing | Zhong et al. (2020) |
| TTA | flip average |
| Ranking | , |
4.3 Metrics
We report quadratic weighted kappa (QWK) Cohen (1968), mean squared error (MSE), accuracy (ACC), macro-F1, and recall. For QWK/ACC/F1/Recall, continuous predictions are rounded to nearest integer grade and clipped to .
5 Results
5.1 Main comparison
Table 2 reports that AGE-Net achieves the best overall performance (QWK 0.90170.0045; MSE 0.23490.0028) compared with strong CNN baselines and representative transformer baselines under our current setting.
| Method | QWK | MSE | ACC | F1 | Rec |
| AGE-Net (Full) | 0.90170.0045 | 0.23490.0028 | 0.78390.0056 | 0.80120.0053 | 0.79190.0060 |
| AGE-Net (w/o Rank) | 0.89630.0033 | 0.26000.0083 | 0.77380.0056 | 0.78930.0067 | 0.77970.0056 |
| AGE-Net (w/o AGR) | 0.89320.0018 | 0.26650.0017 | 0.76820.0107 | 0.77860.0131 | 0.76340.0149 |
| AGE-Net (Base ConvNeXt) | 0.89090.0113 | 0.27950.0287 | 0.77010.0051 | 0.77740.0273 | 0.76320.0373 |
| VGG16 Simonyan and Zisserman (2015) | 0.86540.0069 | 0.30460.0102 | 0.71540.0089 | 0.71800.0170 | 0.69550.0165 |
| EfficientNet Tan and Le (2019) | 0.86380.0012 | 0.30780.0060 | 0.70870.0049 | 0.71660.0050 | 0.69390.0058 |
| ConvNeXt-Base Liu et al. (2022) | 0.86020.0077 | 0.32480.0118 | 0.71320.0087 | 0.72950.0019 | 0.72980.0039 |
| Inception-V3 Szegedy et al. (2016) | 0.84750.0075 | 0.32520.0069 | 0.67520.0030 | 0.68550.0093 | 0.65900.0113 |
| DenseNet121 Huang et al. (2017) | 0.84270.0088 | 0.32740.0032 | 0.66290.0126 | 0.60100.0191 | 0.57500.0189 |
| ResNet50 He et al. (2016) | 0.82710.0012 | 0.37110.0021 | 0.62720.0033 | 0.59320.0251 | 0.56290.0195 |
| Swin-T Liu et al. (2021) | 0.56030.0329 | 0.77670.0493 | 0.35310.0140 | 0.29350.0179 | 0.33630.0101 |
| ViT-B Dosovitskiy et al. (2021) | 0.34730.0298 | 1.04130.0084 | 0.26820.0168 | 0.16120.0217 | 0.25190.0112 |
5.2 Ablation study
Table 3 isolates the effect of major components. Removing ranking decreases QWK, consistent with the importance of explicit ordinal constraints. Removing AGR leads to a larger drop, highlighting the value of anatomy-aware long-range reasoning.
| Variant | QWK | MSE |
| Full | 0.90170.0045 | 0.23490.0028 |
| w/o Rank | 0.89630.0033 | 0.26000.0083 |
| w/o AGR | 0.89320.0018 | 0.26650.0017 |
| Base ConvNeXt | 0.89090.0113 | 0.27950.0287 |
| (TBD) w/o SSF | – | – |
| (TBD) w/o DFR | – | – |
| (TBD) w/o COE (heteroscedastic) | – | – |
5.3 Per-grade analysis (planned / in-progress)
We will add per-grade precision/recall and confusion matrices to better characterize borderline failure modes (e.g., confusion between grades 1/2/3). These analyses often reveal whether improvements concentrate on adjacent grade disambiguation.
5.4 Uncertainty quality (planned / in-progress)
To validate uncertainty, we will evaluate: (i) binned uncertainty vs. absolute error; (ii) selective prediction via risk–coverage curves; (iii) correlation metrics (Spearman/Pearson) between uncertainty and error.
5.5 Robustness under perturbations (planned / in-progress)
We will evaluate robustness under brightness/contrast shifts, Gaussian noise, blur, and occlusion. We will report QWK/MSE degradation curves and compare baseline vs. AGE-Net to quantify resilience benefits from SSF/DFR/AGR.
6 Discussion
Automated KL grading from knee radiographs is challenging because clinically meaningful evidence is often subtle, spatially heterogeneous, and intrinsically ambiguous near adjacent-grade boundaries. AGE-Net is designed around these characteristics. The observed performance gains are consistent with the complementary effects of its components: SSF encourages frequency-aware enhancement of weak texture and contrast cues while retaining spatial selectivity through a lightweight gate, which is desirable when early degenerative findings are low signal-to-noise. AGR explicitly aggregates information across anatomically related regions by message passing over pooled tokens with adaptive neighborhoods, supporting non-local integration of compartmental and alignment context that may be difficult to capture with strictly local operators alone. DFR emphasizes boundary-sensitive feature transitions, a mechanism conceptually aligned with radiographic hallmarks such as joint-space narrowing and contour changes, while its residual formulation preserves the underlying representation and limits overfitting to spurious edges. Collectively, these design choices promote multi-scale evidence integration—local appearance, global anatomical context, and boundary-related structure—that mirrors how OA severity is assessed in practice.
The learning formulation further targets clinically plausible behavior. Treating KL as ordinal rather than as five independent classes and adding a pairwise ranking constraint reduces monotonicity violations and discourages degenerate solutions that can achieve low pointwise error yet produce implausible relative ordering across studies. The evidential COE-Head outputs a severity estimate together with a deterministic uncertainty signal, which may be useful for downstream selective prediction (e.g., deferring high-uncertainty cases for secondary review) in settings where borderline studies and inter-reader variability are common. Several considerations are important for translation. First, uncertainty should be evaluated with task-relevant quality measures, including calibration and selective prediction behavior (risk–coverage or coverage–accuracy curves), and should be interpreted at the level of clinically actionable decisions rather than as an abstract scalar. Second, generalizability should be established through external validation across acquisition sites, scanners, and patient populations, with subgroup analyses to identify potential performance disparities. Third, because backbone choice can interact with pretraining and optimization in medical imaging, comparisons with attention-based alternatives should be conducted under controlled conditions with matched pretraining and tuning to isolate architectural contributions. Addressing these points would strengthen the evidence base for uncertainty-aware and ordinal-consistent KL grading in real-world radiographic workflows.
7 Conclusion
We presented AGE-Net, an ordinal and uncertainty-aware framework for automated Kellgren–Lawrence (KL) grading from knee radiographs. AGE-Net addresses key challenges of radiographic OA assessment—subtle appearance cues, anatomically distributed evidence, and ambiguity near adjacent-grade boundaries—by integrating three complementary enhancement modules: Spectral–Spatial Fusion (SSF) for frequency-aware texture and contrast enhancement, Anatomical Graph Reasoning (AGR) for anatomy-aligned non-local dependency modeling, and Differential Refinement (DFR) for boundary-sensitive feature refinement. To better reflect the ordinal nature of KL grading and to support uncertainty-aware decision making, AGE-Net employs an evidential Normal–Inverse-Gamma (NIG) regression head and a pairwise ordinal ranking constraint, yielding predictions that are both severity-consistent and accompanied by an uncertainty signal.
Under our evaluation protocol with three random seeds, AGE-Net achieved a quadratic weighted kappa of and a mean squared error of , with ablation results indicating that the proposed modules provide complementary and consistent improvements over a strong ConvNeXt baseline. These findings suggest that combining frequency-aware enhancement, structured non-local reasoning, and boundary-focused refinement is a practical strategy for improving KL grading from radiographs. Future work will focus on rigorous characterization of uncertainty quality (e.g., calibration and selective prediction behavior), robustness under distribution shifts, and clinically interpretable model explanations, as well as external validation to assess generalizability across acquisition settings and patient populations.
Appendix A Additional implementation notes
A.1 Training details and stability
We used automatic mixed precision to improve training throughput and reduce memory consumption. To stabilize evaluation and reduce epoch-to-epoch variance, we maintained an exponential moving average (EMA) of model parameters and reported results using the EMA weights at test time. Because the backbone was initialized from pretrained ConvNeXt weights, we adopted layer-wise learning-rate decoupling: the backbone was optimized with a smaller learning rate to mitigate catastrophic forgetting of transferable representations, whereas the newly introduced modules and prediction head used a larger learning rate to learn task-specific mappings. Unless otherwise specified, all ablations followed the same optimization schedule to ensure fair comparison.
We also found it important to handle interactions between augmentation and ordinal constraints. In particular, we disabled the ranking loss when Mixup augmentation was active because Mixup produces interpolated (soft) labels, making the strict inequality ill-defined and potentially introducing inconsistent supervision in the ordinal pair set. In such cases, the model was trained using the evidential regression objective alone.
A.2 Test-time augmentation
At inference, we applied a lightweight flip-based test-time augmentation (TTA) by averaging predictions from the original image and its horizontally flipped counterpart. This procedure can reduce prediction variance and improve robustness to minor pose and acquisition differences without increasing training complexity. For the evidential head, we averaged the predicted severity scores and used the corresponding averaged score for final grade rounding; uncertainty-related parameters were handled consistently with the score aggregation used in our implementation.
Appendix B Extended ablations and statistical analysis
B.1 Planned extended ablations
To further characterize the contribution of each design choice and assess robustness of hyperparameter settings, we plan the following additional ablations:
w/o SSF: remove frequency modulation and spatial gating to isolate the contribution of spectral–spatial enhancement.
w/o DFR: remove differential refinement to isolate boundary-sensitive enhancement.
w/o COE: replace the evidential NIG head with a conventional heteroscedastic regression head and compare predictive uncertainty behavior.
sensitivity: vary in AGR (e.g., 5/9/13) to study the effect of neighborhood size on non-local reasoning.
Token grid size: vary the pooling grid resolution (e.g., , , ) to evaluate the trade-off between token granularity and computational cost.
B.2 Planned statistical analysis
To quantify the reliability of observed improvements, we plan to report paired comparisons across random seeds when evaluating AGE-Net against major ablations and the ConvNeXt baseline. Specifically, we will apply paired statistical tests on per-seed summary metrics (e.g., paired -tests on QWK) and report corresponding effect sizes. Where appropriate, we will additionally compute confidence intervals (e.g., via bootstrap resampling) to complement -values and better convey the magnitude and uncertainty of performance differences.